OpenStudy (anonymous):

Half-life is the time it takes for half of a sample of a radioactive element to decay. The function M(t)=P(1/2)^t/h can be used to calculate the mass remaining if the half-life is h and the initial mass is P. The half-life of radium is 1620 years. a) How many years will it take until the laboratory has only 4 g of radium?

OpenStudy (anonymous):

This is what I have so far: M(t) = P(1/2)^t/1620 4 = 5(1/2)^t/1620 (1/2)^t/1620 = 4/5 log1/2(4/5)=t/1620 1620log1/2(4/5)=t and then I'm stuck.......

OpenStudy (anonymous):

log1/2(4/5) = (log4/5)/(log1/2)